# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ Abstract.
Charney and Morris-Wright showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes and the actions on them. The authors developed their study and clarified when acylindrical hyperbolicity holds for Artin groups of infinite type associated with graphs that are not cones. In this paper, we introduce reduced clique-cube complexes. By using them, we show acylindrical hyperbolicity of irreducible Artin groups associated with graphs that are not cliques, that is, irreducible Artin groups that are not free of infinity. As an application, we see that the centers of such Artin groups are finite, and that actually they are trivial in many cases. Also as a corollary, we see that irreducible Artin groups of infinite type and of type FC are acylindrically hyperbolic.
@@ 1. Introduction
Artin groups, also called Artin-Tits groups, have been widely studied since their introduction by Tits [36]. Let $\Gamma$ be a finite simple graph with the vertex set $V(\Gamma)$ and the edge set $E(\Gamma)$. Each edge $e$ has two end vertices, which we denote as $s_{e}$ and $t_{e}$. We suppose that any edge $e$ is labeled by an integer $\mu(e)\geq 2$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the following presentation
@@
Here, $\Gamma$ is called a defining graph of $A_{\Gamma}$ and the rank of $A_{\Gamma}$ is defined as #$\#V(\Gamma)$. Free abelian groups, free groups, and braid groups are typical examples of Artin groups. For $U\subset V(\Gamma)$, we denote the subgroup generated by $U$ as $A_{U}$, and call such a subgroup a standard parabolic subgroup. Adding the relation $v^{2}=1$ for all $v\in V(\Gamma)$ to (1.1) produces the associated Coxeter group $W_{\Gamma}$. Let $\Gamma^{t}$ be the Coxeter graph defined from the presentation of $W_{\Gamma}$. That is, $\Gamma^{t}$ is a finite simple graph with edge labels in $\mathbb{Z}_{\geq 2}\cup\{\infty\}$, which is gained from $\Gamma$ by deleting all edges with label $2$ and adding an edge with label $\infty$ between every pair of vertices that are not connected by an edge in $\Gamma$. Although $\Gamma^{t}$ is used in research related to Coxeter groups and traditional treatments of Artin groups, we mainly use $\Gamma$ in accordance with many recent studies on Artin groups in geometric group theory. We use the associated Coxeter graph $\Gamma^{t}$ as an aid in this paper. See Figures 1 and 2.
Figure 1. The defining graphs $\Gamma=\Gamma_{0}\ast\Gamma_{1}$ and the associated Coxeter graphs $\Gamma^{t}$. In each $\Gamma$, $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ is an indecomposable factor with two vertices.
Figure 2. The defining graphs $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$ and the associated Coxeter graphs $\Gamma^{t}$. In each $\Gamma$, $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ and $\Gamma_{2}$ are indecomposable factors with two vertices.
In terms of the properties of $W_{\Gamma}$ or $\Gamma^{t}$, we can define several important classes of Artin groups. The Artin group $A_{\Gamma}$ is said to be of spherical type if $W_{\Gamma}$ is finite. Otherwise, it is said to be of infinite type. The Artin group $A_{\Gamma}$ is said to be irreducible if $W_{\Gamma}$ is irreducible, that is, $\Gamma^{t}$ is connected. If $\Gamma^{t}$ does not contain edges labeled by $\infty$, then we say $A_{\Gamma}$ is free of infinity. By definition, $A_{\Gamma}$ is free of infinity if and only if $\Gamma$ is a clique. In this paper, we focus on Artin groups which are not free of infinity, that is, Artin groups associated with graphs that are not cliques.
In general, compared with Artin groups, Coxeter groups are relatively well understood. For example, it is well-known that an infinite Coxeter group $W_{\Gamma}$ is irreducible if and only if $W_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups ([32] and [36]). However, it is unclear whether $A_{\Gamma}$ is irreducible if and only if $A_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups. We remark that many basic questions for Artin groups remain open (refer to [12] and [20]).
In order to study Artin groups, it is often effective to investigate hidden nonpositively curved or negatively curved properties of them. In this paper, we consider the following problem [23, Conjecture B].
@@
Acylindrical hyperbolicity is an important property in geometric group theory. We can find various applications of acylindrical hyperbolicity in [16], [33], and [34] for example.
We note that reducible Artin groups are not acylindrically hyperbolic, since acylindrically hyperbolic groups cannot be directly decomposed into two infinite subgroups ([33, Corollary 7.3]). We also note that Artin groups with infinite centers are not acylindrically hyperbolic, since acylindrically hyperbolic groups do not permit infinite centers ([33, Corollary 7.3]). Although irreducible Artin groups of spherical type have infinite cyclic center ([4] and [17]) and thus are not acylindrically hyperbolic, the central quotients for them of rank at least $2$ are acylindrically hyperbolic (see [2], [24], [5] for braid groups, and [10] for the general case).
Many affirmative partial answers for Problem 1.1 are known. Indeed, the following irreducible Artin groups of infinite type are known to be acylindrically hyperbolic:
@@
  • Two-dimensional Artin groups, that is, Artin groups such that every triangle with three vertices $v_{1},v_{2},v_{3}$ of the defining graphs satisfies $\frac{1}{\mu((v_{1},v_{2}))}+\frac{1}{\mu((v_{2},v_{3}))}+\frac{1}{\mu((v_{3},v_{1}))}\leq 1$ ([42]).
  • Artin groups associated with graphs that are not joins ([14]).
  • Artin groups associated with graphs that are not cones ([29]).
  • Relatively extra-large-type Artin groups with mild conditions ([21]).
  • Even Artin groups of type FC and Artin group of type FC that visually splits over a standard parabolic subgroup of spherical type ([13]).
As a general strategy for Problem 1.1, we consider the case that the defining graph is a clique and the case that the defining graph is not a clique separately. Note that Artin groups in the latter case are automatically of infinite type. In this paper we give the affirmative answer for Problem 1.1 for the latter case. In fact, our main theorem is the following:
@@
  1. (1)
    $A_{\Gamma}$ is irreducible;
  2. (2)
  3. (3)
    $A_{\Gamma}$ is acylindrically hyperbolic;
  4. (4)
    $A_{\Gamma}$ is virtually directly indecomposable, that is, it cannot be decomposed as a direct product of two infinite subgroups.
From Theorem 1.2, we find that many irreducible Artin groups of infinite type are acylindrically hyperbolic, e.g., the Artin groups associated with the defining graphs in Figures 1 and 2. We also apply Theorem 1.2 to Artin groups of infinite type and of type FC.
@@
Note that, when the defining graph of type FC is a clique, the Artin group is of spherical type.
We have an application to the so-called center conjecture, which claims that the center of any irreducible Artin group of infinite type is trivial.
@@
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then the center $Z(A_{\Gamma})$ is finite. In addition, $Z(A_{\Gamma})$ is contained in $Z(A_{\Gamma_{0}})$ where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$. In particular, if $Z(A_{\Gamma_{0}})$ is trivial or torsion-free, then $Z(A_{\Gamma})$ is trivial.
@@
Also we have a corollary for direct products.
@@
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Then if $A_{\Gamma_{0}}$ is directly indecomposable or no nontrivial finite subgroup which is a factor of a direct decompositon, then $A_{\Gamma}$ is directly indecomposable.
@@
Let $A_{\Gamma}$ be an irreduicble Artin group associated with $\Gamma$ that is not a clique. Suppose that $\Gamma_{0}$ consists of three vertices at most. See Figures 1 and 2. Then $Z(A_{\Gamma_{0}})$ is known to be trivial or torsion-free, and thus $Z(A_{\Gamma})$ is trivial by Corollary 1.4. Also $A_{\Gamma_{0}}$ is directly indecomposable or no nontrivial finite subgroup which is a factor of a direct decompositon, and thus $A_{\Gamma}$ is directly indecomposable by Corollary 1.5.
@@
Goldsborough and Vaskou, who are the authors of [22], pointed us out that results in [22] and Theorem 1.2 induce acylindrical hyperbolicity of random Artin groups. The following is based on their remark. By referring to [22] for some definitions, we can write down the claim precisely as follows: an Artin group picked at random relatively to any non-decreasing divergent function $f:\mathbb{N}\to\mathbb{N}$ is asymptotically almost surely acylindrically hyperbolic. Indeed the case of $f(N)\succ N^{\frac{3}{2}}$ and the case of $f(N)\prec N^{1-\alpha}$ for some $10<\alpha<$ 1 are covered in Corollaries 3.4 and 4.3 in [22], respectively. The missing case was unknown (see Question 1.6 in [22]), but now it is resolved as follows: we consider the case of $f(N)\prec N^{2}$, which contains the missing case. By Lemma 2.7, Theorem 2.9 (2) and Lemma 2.10 in [22], an Artin group picked at random relatively to $f$ asymptotically almost surely is irreducible and not free of infinity. Thus it is acylindrically hyperbolic by Theorem 1.2.
@@
In this paper, for the proof of Theorem 1.2, we develop the argument in Charney and Morris-Wright [14] and the authors [29]. Charney and Morris-Wright [14] showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes, which are CAT(0) cube complexes, and the isometric actions on them. They constructed a WPD contracting element of such an Artin group with respect to the isometric action on the clique-cube complex. The authors [29] developed their study and treated Artin groups associated with graphs that are not cones. To treat the case of Artin groups associated with graphs that are not cliques, we need further development and more delicate arguments. Indeed, we introduce the reduced clique-cube complex. It is a convex subcomplex of the clique-cube complex, and works more appopriately in our case than the clique-cube complex. Combining with careful discussions on the local geometry of the cube complexes, we construct a candidate of a WPD contracting element with respect to the action on the reduced clique-cube complex. We also apply some algebraic results on standard parabolic subgroups of Artin groups when we show that the candidate of a WPD contracting element is actually WPD contracting.
@@
The remainder of this paper is organized as follows. Section 2 contains some preliminaries. Section 2.1 is for preliminaries regarding acylindrically hyperbolic groups, WPD contracting elements, and CAT(0)-cube complexes. Section 2.2 presents preliminaries on defining graphs of Artin groups. Section 2.3 presents facts on standard parabolic subgroups of Artin groups and prove a lemma on intersections of standard parabolic subgroups and their conjugates in Artin groups. Section 2.4 treats clique-cube complexes. Section 2.5 deals with the definition of joins of graphs and some properties of them. In Section 3, we introduce reduced clique-cube complexes and study the local geometry. In Section 4, we give a proof of Theorem 1.2. Our main task is to construct a candidate WPD contracting element and show that it really is a WPD contracting element. Also we show Corollaries 1.4 and 1.5.
@@
2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0)-cube complexes
We collect some definitions and properties related to acylindrical hyperbolicity, WPD contracting elements, and CAT(0)-cube complexes that will be used later in the paper. See [18] and the references therein for details.
@@ 2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0)-cube complexes
  • $\gamma$ is WPD if, for every $D\geq 0$ and $x\in X$, there exists some $M\geq 1$ such that #$\#\{g\in G|\ d_{X}(x,g(x)),d_{X}(\gamma^{M}(x),g\gamma^{M}(x))\leq D\}<\infty$;
  • $\gamma$ is contracting if $\gamma$ is loxodromic, that is, there exists $x_{0}\in X$ such that $\mathbb{Z}\to X;n\mapsto\gamma^{n}(x_{0})$ is a quasi-isometry onto the image $\gamma^{\mathbb{Z}}x_{0}:=\{\gamma^{n}(x_{0})|n\in\mathbb{Z}\}$, and $\gamma^{\mathbb{Z}}x_{0}$ is contracting, that is, there exists $B\geq 0$ such that the diameter of the nearest-point projection of any ball that is disjoint from $\gamma^{\mathbb{Z}}x_{0}$ onto $\gamma^{\mathbb{Z}}x_{0}$ is bounded by $B$.
@@
CAT(0) cube complexes are considered as generalized trees in higher dimensions. The following is a precise definition (see [7, p.111]).
@@
Let $X$ be a CAT(0) cube complex. We define an equivalence relation for the edges of $X$ as the transitive closure of the relation identifying two parallel edges of a square. For an equivalence class, a hyperplane is defined as the union of the midcubes transverse to the edges belonging to the equivalence class. Then, for any edge belonging to the equivalence class, the hyperplane is said to be dual to the edge.
For a hyperplane $J$, we denote the union of the cubes intersecting $J$ as $N(J)$, that is, the smallest subcomplex of $X$ containing $J$. We denote the union of the cubes not intersecting $J$ as $X\!\setminus\!\!\setminus J$, that is, the largest subcomplex of $X$ not intersecting $J$.
@@
Let $X$ be a CAT(0) cube complex and $J$ be a hyperplane. Then, $X\!\setminus\!\!\setminus J$ has exactly two connected components.
The two connected components of $X\!\setminus\!\!\setminus J$ are often denoted as $J^{+}$ and $J^{-}$, respectively.
The following is given in [29, Definition 2.7].
@@
The following is part of [18, Theorem 3.3], and is used in the proof of Theorem 1.2.
@@
    $J$ and $J^{\prime}$ are strongly separated, that is, no hyperplane can intersect both $J$ and $J^{\prime}$;
@@
    $\gamma$ skewers $J$ and $J^{\prime}$, that is, we have connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$ and $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$ such that $\gamma^{n}(J^{+})\subsetneq J^{\prime+}\subsetneq J^{+}$ for some $n\in\mathbb{N}$;

### v3

@@ Abstract.
Charney and Morris-Wright showed acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes and the actions on them. The authors developed their study and clarified when acylindrical hyperbolicity holds for Artin groups of infinite type associated with graphs that are not cones. In this paper, we introduce reduced clique-cube complexes. Using these complexes, we show acylindrical hyperbolicity of irreducible Artin groups associated with graphs that are not cliques, i.e., irreducible Artin groups that are not free of infinity. One application of our main result is that the centers of such Artin groups are finite, and that actually they are trivial in many cases. Another corollary is that irreducible Artin groups of infinite type and of type FC are acylindrically hyperbolic.
@@ 1. Introduction
Artin groups, also called Artin-Tits groups, have been widely studied since their introduction by Tits [36]. Let $\Gamma$ be a finite simple graph with the vertex set $V(\Gamma)$ and the edge set $E(\Gamma)$. Each edge $e$ has two endpoints, which we denote by $s_{e}$ and $t_{e}$. We suppose that any edge $e$ is labeled by an integer $\mu(e)\geq 2$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the following presentation:
@@
[…]
Figure 1. An example of a defining graph $\Gamma$ and the associated Coxeter graph $\Gamma^{t}$. Thick solid lines indicate edges labeled by $3$, and dotted lines indicate edges labeled by $2$. In $\Gamma^{t}$, thin solid lines denote edges labeled by $\infty$.
In terms of the properties of $W_{\Gamma}$ or $\Gamma^{t}$, we can define several important classes of Artin groups. The Artin group $A_{\Gamma}$ is said to be of spherical type if $W_{\Gamma}$ is finite. Otherwise, it is said to be of infinite type. The Artin group $A_{\Gamma}$ is said to be irreducible if $W_{\Gamma}$ is irreducible; that is, when $\Gamma^{t}$ is connected. If $\Gamma^{t}$ does not contain edges labeled by $\infty$, then we say $A_{\Gamma}$ is free of infinity. By definition, $A_{\Gamma}$ is free of infinity if and only if $\Gamma$ is a clique. In this paper, we focus on Artin groups that are not free of infinity; that is, Artin groups associated with graphs that are not cliques.
In general, compared with Artin groups, Coxeter groups are relatively well understood. For example, it is well known that an infinite Coxeter group $W_{\Gamma}$ is irreducible if and only if $W_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups ([32] and [36]). However, it is unclear whether $A_{\Gamma}$ is irreducible if and only if $A_{\Gamma}$ cannot be directly decomposed into two nontrivial subgroups. We remark that many basic questions about Artin groups remain open (refer to [12] and [20]).
To study Artin groups, it is often effective to investigate their hidden nonpositively curved or negatively curved properties. In this paper, we consider the following problem [23, Conjecture B].
@@
Acylindrical hyperbolicity is an important property in geometric group theory. We can find various applications of acylindrical hyperbolicity in [16], [33] and [34] for example.
It is known that reducible Artin groups are not acylindrically hyperbolic, since acylindrically hyperbolic groups cannot be directly decomposed into two infinite subgroups ([33, Corollary 7.3]). We also note that Artin groups with infinite centers are not acylindrically hyperbolic, since acylindrically hyperbolic groups do not admit infinite centers ([33, Corollary 7.3]). Although irreducible Artin groups of spherical type have infinite cyclic center ([4] and [17]) and thus are not acylindrically hyperbolic, their central quotients are acylindrically hyperbolic when the rank is at least $2$ (see [2], [24], [5] for braid groups and [10] for the general case).
Many affirmative partial answers to Problem 1.1 are known. Indeed, the following irreducible Artin groups of infinite type are known to be acylindrically hyperbolic:
@@
  • Two-dimensional Artin groups; that is, Artin groups for which every triangle with three edges $e_{1},e_{2},e_{3}$ in the defining graphs satisfy $\frac{1}{\mu(e_{1})}+\frac{1}{\mu(e_{2})}+\frac{1}{\mu(e_{3})}\leq 1$ ([42]).
  • Artin groups of which defining graphs that are not joins ([14]).
  • Artin groups of which defining graphs that are not cones ([29]).
  • Relatively extra-large-type Artin groups under mild conditions ([21]).
  • Even Artin groups of type FC, as well as Artin groups of type FC that visually split over standard parabolic subgroups of spherical type ([13]).
As a general strategy for Problem 1.1, we separately consider the case in which the defining graph is a clique and the case in which the defining graph is not a clique. Note that Artin groups in the latter case are automatically of infinite type. In this paper, we give an affirmative answer to Problem 1.1 in the latter case. Our main theorem is the following:
@@
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$. Suppose that $\Gamma$ is not a clique. Then $A_{\Gamma}$ is acylindrically hyperbolic.
From Theorem 1.2, we find that many irreducible Artin groups of infinite type are acylindrically hyperbolic, e.g., the Artin group associated with the defining graph in Figure 1. We also apply Theorem 1.2 to Artin groups that are both of infinite type and of type FC.
@@
Note that, an Artin group of type FC is of spherical type if the defining graph is a clique.
Before we give other applications, we define the maximal clique factor of $\Gamma$. Let $\Gamma$ be a simple graph that is not a clique. If a factor of a join decomposition of $\Gamma$ is a clique, then it is called a clique factor. If $\Gamma$ has a clique factor, then it has a unique maximal one, which is called the maximal clique factor and usually denoted by $\Gamma_{0}$ in this paper. See Figures 1 and 2. For convenience, if $\Gamma$ has no nontrivial clique factor, we set $\Gamma_{0}=\emptyset$ and refer to it as the maximal clique factor, as well. See Section 2.5 for details related to join decompositions, and Figures 4 and 5 for other examples.
Figure 2. $\Gamma$ in Figure 1 has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$, and $\Gamma_{0}$ is the maximal clique factor. Later we denote $\Gamma_{\ast}=\Gamma_{1}\ast\Gamma_{2}$. See Section 2.5.
We apply Theorem 1.2 to the so-called center conjecture, which states that the center of any irreducible Artin group of infinite type is trivial.
@@
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then the center $Z(A_{\Gamma})$ is finite. Moreover, $Z(A_{\Gamma})$ is contained in $Z(A_{\Gamma_{0}})$. In particular, if $Z(A_{\Gamma_{0}})$ is trivial or torsion-free, then $Z(A_{\Gamma})$ is trivial. Here, $\Gamma_{0}$ is the maximal clique factor of $\Gamma$.
@@
We also obtain a corollary concerning direct products.
@@
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Then $A_{\Gamma}$ is virtually directly indecomposable; that is, it cannot be decomposed as a direct product of two infinite subgroups. Moreover, if $A_{\Gamma_{0}}$ contains no nontrivial finite subgroup that is a factor of a direct decomposition of $A_{\Gamma_{0}}$, then $A_{\Gamma}$ is directly indecomposable. Here, $\Gamma_{0}$ is the maximal clique factor of $\Gamma$.
@@
Let $A_{\Gamma}$ be an irreducible Artin group associated with $\Gamma$ that is not a clique. Suppose that the maximal clique factor $\Gamma_{0}$ consists of three vertices or fewer. See Figures 1 and 2. Then $Z(A_{\Gamma_{0}})$ is known to be trivial or torsion-free, and thus $Z(A_{\Gamma})$ is trivial by Corollary 1.4. Also, $A_{\Gamma_{0}}$ is known to be torsion-free, and thus $A_{\Gamma}$ is directly indecomposable by Corollary 1.5.
@@
Goldsborough and Vaskou, who are the authors of [22], pointed to us out that the results in [22] and Theorem 1.2 imply acylindrical hyperbolicity of random Artin groups. The following is based on their remark. By referring to [22] for some definitions, we can state the resulting implication precisely: an Artin group picked at random relatively to any non-decreasing divergent function $f:\mathbb{N}\to\mathbb{N}$ is asymptotically almost surely acylindrically hyperbolic. Indeed the case of $f(N)\succ N^{\frac{3}{2}}$ and the case of $f(N)\prec N^{1-\alpha}$ for some $10<\alpha<$ 1 are covered in Corollaries 3.4 and 4.3 in [22], respectively. The missing case was unknown (see Question 1.6 in [22]), but now it is resolved as follows: we consider the case of $f(N)\prec N^{2}$, which contains the missing case. By Lemma 2.7, Theorem 2.9 (2) and Lemma 2.10 in [22], an Artin group picked at random relatively to $f$ asymptotically almost surely is irreducible and not free of infinity. Thus, it is acylindrically hyperbolic by Theorem 1.2.
@@
In this paper, for the proof of Theorem 1.2, we develop arguments from [14] and [29]. Charney and Morris-Wright [14] established the acylindrical hyperbolicity of Artin groups of infinite type associated with graphs that are not joins, by studying clique-cube complexes, which are CAT(0) cube complexes, and the isometric actions on them. They constructed a WPD contracting element of such an Artin group with respect to the isometric action on the clique-cube complex. We note that the existence of a WPD contracting element is a sufficient condition for acylindrical hyperbolicity [1]. The authors [29] developed this approach and treated Artin groups associated with graphs that are not cones. To treat the case of Artin groups associated with graphs that are not cliques, we need additional development and more refined arguments. Indeed, we introduce the reduced clique-cube complex. It is a convex subcomplex of the clique-cube complex, and works more appropriately in our case than the clique-cube complex. Together with careful analysis of the local geometry of the cube complexes, we construct a candidate of a WPD contracting element with respect to the action on the reduced clique-cube complex. We also apply some algebraic results on standard parabolic subgroups of Artin groups to show that the candidate is actually a WPD contracting element.
@@
The remainder of this paper is organized as follows. Section 2 contains some preliminaries. Section 2.1 is for preliminaries regarding acylindrically hyperbolic groups, WPD contracting elements, and CAT(0) cube complexes. Section 2.2 presents preliminaries on defining graphs of Artin groups. Section 2.3 presents facts on standard parabolic subgroups of Artin groups and gives a lemma on intersections of standard parabolic subgroups and their conjugates in Artin groups. Section 2.4 treats clique-cube complexes. Section 2.5 deals with the definition of joins of graphs and some related properties. In Section 2.6, we prepare some operations on matrices over a group, which are used when we construct a WPD contracting element in Section 4. In Section 3, we introduce reduced clique-cube complexes and study the local geometry. In Section 4, we state Theorem 4.1 which contains Theorem 1.2 and adds information, and prove it. Our main task is to construct a candidate WPD contracting element and show that it is really a WPD contracting element. We also prove Corollaries 1.4 and 1.5. In Section 5, we treat an example to aid understanding of Section 4.
@@
2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0) cube complexes
We collect some definitions and properties related to acylindrical hyperbolicity, WPD contracting elements, and CAT(0) cube complexes for later use. See [18] and the references therein for details.
@@ 2.1. Acylindrical hyperbolicity, WPD contracting elements, and CAT(0) cube complexes
  • $\gamma$ is weakly properly discontinuous (WPD) if, for every $D\geq 0$ and $x\in X$, there exists some $M\geq 1$ such that
    #$\#\{g\in G|\ d_{X}(x,g(x)),d_{X}(\gamma^{M}(x),g\gamma^{M}(x))\leq D\}<\infty;$
  • $\gamma$ is contracting if $\gamma$ is loxodromic, that is, there exists $x_{0}\in X$ such that the map $\mathbb{Z}\to X;n\mapsto\gamma^{n}(x_{0})$ is a quasi-isometry onto the image $\gamma^{\mathbb{Z}}x_{0}:=\{\gamma^{n}(x_{0})|n\in\mathbb{Z}\}$, and $\gamma^{\mathbb{Z}}x_{0}$ is contracting; that is, there exists $B\geq 0$ such that the diameter of the nearest-point projection of any ball that is disjoint from $\gamma^{\mathbb{Z}}x_{0}$ onto $\gamma^{\mathbb{Z}}x_{0}$ is bounded by $B$.
@@
CAT(0) cube complexes are regarded as generalized trees in higher dimensions. The following is a precise definition (see [7, p.111]).
@@
Let $X$ be a CAT(0) cube complex. We define an equivalence relation on the edges of $X$ as the transitive closure of the relation identifying two parallel edges of a square. For an equivalence class, a hyperplane is defined as the union of the midcubes transverse to the edges belonging to the equivalence class. Then, for any edge belonging to the equivalence class, the hyperplane is said to be dual to the edge.
For a hyperplane $J$, we denote the union of the cubes intersecting $J$ as $N(J)$; that is, the smallest subcomplex of $X$ containing $J$. We denote the union of the cubes not intersecting $J$ as $X\!\setminus\!\!\setminus J$; that is, the largest subcomplex of $X$ not intersecting $J$.
@@
Let $X$ be a CAT(0) cube complex and $J$ a hyperplane. Then, $X\!\setminus\!\!\setminus J$ has exactly two connected components.
The two connected components of $X\!\setminus\!\!\setminus J$ are often denoted by $J^{+}$ and $J^{-}$, respectively.
The following is adapted from [29, Definition 2.7].
@@
The following is based on part of [18, Theorem 3.3] and is used in the proof of Theorem 1.2.
@@
    $J$ and $J^{\prime}$ are strongly separated; that is, no hyperplane intersects both $J$ and $J^{\prime}$;
@@
    $\gamma$ skewers $J$ and $J^{\prime}$; that is, we have connected components $J^{+}$ of $X\!\setminus\!\!\setminus J$ and $J^{\prime+}$ of $X\!\setminus\!\!\setminus J^{\prime}$ such that $\gamma^{n}(J^{+})\subsetneq J^{\prime+}\subsetneq J^{+}$ for some $n\in\mathbb{N}$;